Grade 12Physical Chemistry

In following first order reaction A+reagent give product and B+reagent give product,tge ratio of K1 to K2 if only 50percent of B and 94 percent of A hsd reacted

Aniket
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
For A+,
K1 = ln(A/0.06A).........(1)
For B+
K2 = ln(B/0.5B)........(2)
Dividing both the eqns. we get the ratio as:
K1/K2 = 1*0.5/0.06 => 50/6 => 25/3 or 25: 3